10g bullet is fired from a 5 kg gun with a velocity of 400 m/s.What is the speed of the recoil.


Is the answer -- 0.8m/s


If not then provide the correct solution

Grade:11

4 Answers

suryakanth AskiitiansExpert-IITB
13 years ago

Dear nitin,

Your answer is correct

Vg = - (mb/mg)* Vb

    = - ((10* 10^(-3))/5) * 400

    = - (4/5)

    = - 0.8 m/s

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Suryakanth –IITB

aditya suman
13 years ago

the answer 0.8m/s is correct.

shreyas ramoji
13 years ago

it is correct, 

according to law of conservation of momentum,

m1v1=m2v2

v2 = m1v1/m2

v2 = (10x10-3)(400)/5

v2 = (0.01x400)/5

v2 = 4/5 = 0.8 ms-1

 


vikas askiitian expert
13 years ago

your ans is corect

 from conservation of momentam

  M1.V1 + mv2 =0

              V1=-mv2/M

                  =-4/5=-0.8m/s